OpenStudy (unklerhaukus):

Can someone please check this; \[\begin{equation*} f(t)=\cos(t)\big(h(t-\pi)-h(t)\big) \end{equation*}\] \[ \newcommand\dd[1]{\,\mathrm d#1} % infinitesimal \newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd #4}} % integral _{a}^{b}{f(x)}\dd x \begin{align*} \intl0\infty{f(t)e^{-pt}}t \end{align*} \]

OpenStudy (unklerhaukus):

\[\newcommand\dd[1]{\,\mathrm d#1} % infinitesimal \newcommand\intl[4]{\int\limits_{#1}^{#2}{#3}{\dd #4}} % integral _{a}^{b}{f(x)}\dd x \begin{align*} \intl0\infty{f(t)e^{-pt}}t&=\intl0\infty{\cos(t)\big(h(t-\pi)-h(t)\big)e^{-pt}}t\\ &=\intl0\infty{\cos(t)h(t-\pi)e^{-pt}}t-\intl0\infty{\cos(t)h(t)e^{-pt}}t\\ &=\intl\pi\infty{\cos(t)e^{-pt}}t-\intl0\infty{\cos(t)e^{-pt}}t\\ &=\left.\frac{e^{-pt}\big(-p\cos t+\sin t\big)}{(-p)^2+1^2}\right|_\pi^\infty-\left.\frac{e^{-pt}\big(-p\cos t+\sin t\big)}{(-p)^2+1^2}\right|_0^\infty\\ &=\frac{-e^{-p\pi}\big(-p\cos \pi+\sin \pi\big)}{p^2+1}+\frac{\big(-p\cos 0+\sin 0\big)}{p^2+1}\\ &=\frac{-pe^{-p\pi}}{p^2+1}-\frac{p}{p^2+1}\\ \\ &=\frac{-p\big(1+e^{-p\pi }\big)}{p^2+1}\\ \end{align*}\]

OpenStudy (whpalmer4):

I haven't had my coffee yet, but I don't quite see how you managed to make h(t) vanish

OpenStudy (phi):

I think h is the step function

OpenStudy (whpalmer4):

ah, okay, sure, I didn't realize that we could assume anything about h(t)

OpenStudy (experimentx):

isn't that Laplace transform? .. let p = s

OpenStudy (unklerhaukus):

The unit heaviside step function \[h(t-a) =\begin{cases}0&t<a\\ 1&t≥a\end{cases} \]

OpenStudy (experimentx):